OpenStudy (anonymous):

Solve the inequality -3x + 4 <= x + 8 and show all of your work. You can use <= to represent the less than or equal to symbol.

OpenStudy (anonymous):

-4x<=4

OpenStudy (anonymous):

-3x + 4 <= x +8 -3x <= x + 4 -4x <= 4 x >= -1

OpenStudy (anonymous):

-x<=1

OpenStudy (anonymous):

x>=-1